# Bayesian neural networks with interpretable priors from Mercer kernels

### v1

@@ 3 The Mercer prior
Proposition 3.1 provides an unbiased estimator of the Mercer prior as given by eq. (3.4). The estimate exploits the available spectral representation of the covariance operator $S$ for the measure we wish to draw the BNN from. In this way, we avoid ever needing to compute the integrals explicitly, identify the inverse covariance kernel analytically, or invert a large covariance matrix as is typical when sampling from a GP. At any given step, to compute the estimate, we simply need to minibatch points in the domain in order to approximate the integrals and select random eigenvalues/eigenfunctions associated with $S$. The properties of SGLD guarantee convergence to the correct probability distribution even for all batch sizes equal to one, showing the potential for massive parallelization.
In proposition 3.1, we have chosen to minibatch points in $\Omega$ uniformly, which does not allow the domain to be infinite. This choice was made in order to simplify the method and it is not a strict requirement. If one wished to expand the domain to $\mathbb{R}^{d}$, importance sampling can be used to turn the integrals $\langle u_{\theta},\phi_{n}\rangle$ into expectations. Specifically, if we instead sample $x_{\beta}$ and $x_{\gamma}$ from distributions supported on $\mathbb{R}^{d}$, $q(x)$ and $w(x)$, respectively, we approximate each inner product using
$\langle u_{\theta},\phi_{n}\rangle\approx\sum_{\beta=1}^{M_{1}}\frac{u_{\theta}(x_{\beta})\phi_{n}(x_{\beta})}{q(x_{\beta})},\quad$
@@ 3.1 A cruder approximation
$\mathcal{S}_{N,M_{1},M_{2}}=-\frac{1}{2}\frac{K|\Omega|^{2}}{NM_{1}M_{2}}\sum_{\alpha=1}^{N}\frac{1}{\lambda_{n_{\alpha}}}\left\{\sum_{\beta=1}^{M_{1}}u_{\theta}(x_{\beta})\phi_{n_{\alpha}}(x_{\beta})\right\}\left\{\sum_{\gamma=1}^{M_{2}}u_{\theta}(x_{\gamma})\phi_{n_{\alpha}}(x_{\gamma})\right\},$ (3.7)
@@ 3.2 Specifying Gaussian process covariance functions from the ground up
Observe that the method builds sampling approximations of GPs only using the eigenvalues and eigenfunctions of the covariance kernel. This means that the covariance kernel does not need to be explicitly written. The benefit of this is that we may specify the desired GP completely through the eigenvalues and eigenfunctions, which allows us to construct BNNs for specialized tasks. We describe two possible approaches for constructing a desired kernel.
@@
$k(x,x^{\prime})=\sum_{n=1}^{\infty}\lambda_{n}\phi_{n}(x)\phi_{n}(x^{\prime})$ (3.8)
defines a valid covariance kernel for a GP. Note that this also holds if the series is truncated at some finite number of terms. In this way, we may start by choosing the eigenvalues and eigenfunctions so that the GP follows some desired behavior, e.g., smoothness, which is then transferred into the behavior of the BNN. The importance of this is that for GP kernels commonly used in machine learning tasks, the eigenvalues and eigenfunctions are currently not known or not easily computed. As an example, the eigenvalues and eigenfunctions of the squared exponential kernel $k(x,x^{\prime})=\exp(-\|x-x^{\prime}\|^{2}/\ell^{2})$, which is perhaps the most commonly used kernel, are only available under the Gaussian measure $\mathcal{N}(0,\sigma^{2})$, for some variance $\sigma^{2}>0$ [66]. The eigenfunctions are then defined via transformations of the Hermite polynomials. In numerical experiments not covered in this work, we found that we could not include enough of these eigenfunctions in the expansion to build a good approximation of the GP. This was because after roughly $100$ terms, the computations became very unstable, mostly attributed to the fact that the normalization constants became exceptionally large.
@@
    For all fixed $x\in\Omega$, $k(\cdot,x)\in H_{k}$.
@@
    For all fixed $x\in\Omega$, $u(x)=\langle u,k(\cdot,x)\rangle_{H_{k}}$ for each $u\in H_{k}$.
@@
The second method involves damping an existing covariance with a Mercer representation. One specific target application we have in mind for the Mercer prior is in infinite-dimensional Bayesian inverse problems, where one would look to replace the usual prior Gaussian measure with the corresponding BNN for scalability. In such inverse problems, one often works with a ‘Laplace-like’ covariance, see [50]. This is first described by the baseline Gaussian measure $\mathcal{N}(0,\Delta^{-1})$, where $\Delta$ is the Laplacian with Dirichlet boundary conditions. This measure is famously known as the Brownian bridge measure. On $[0,1]$, the kernel of the inverse Laplacian operator $\Delta^{-1}$ has eigenvalues and corresponding eigenfunctions
$\lambda_{n}$ $=\frac{1}{n^{2}\pi^{2}},$
$\phi_{n}(t)$ $=\sqrt{2}\sin(n\pi t).$
@@ 3.3 Computational cost analysis
$\mathcal{S}_{M_{1}}=\sum_{\beta=1}^{M_{1}}u_{\theta}(x_{\beta})\phi_{n_{\alpha}}(x_{\beta}),$
and similarly for $M_{2}$. Each term $u_{\theta}(x)$ requires a forward pass through a neural network, which costs $O(D)$. Evaluating an eigenfunction $\phi_{n_{\alpha}}(x)$ costs $O(1)$, unless the eigenfunction is a special function, which we will ignore. So, this incurs a cost of $O((M_{1}+M_{2})D)=O(MD)$ per eigen-index if we take $M_{1}=M_{2}=M$. Multiplying by $N$ gives the total leading cost $O(NMD)$. To be completely fair, we acknowledge the fact that SGLD tends to mix very slowly, so the unbiased estimate needs to be computed many times in order to generate one independent posterior draw. If we let $T$ denote the number of SGLD steps needed for an independent sample, the total cost is $O(TNMD)$. The mixing rate can be improved by implementing a variant of SGLD, such as preconditioning [57] or cyclical SGLD [65]. Alternatively, the prior can be characterized by other methods which are known to work with unbiased estimates, such as Hamiltonian Monte Carlo with energy conserving subsampling [12], or with variational inference [6].
@@
If we replace the GP with a BNN equipped with the Mercer prior, generating a sample from the posterior predictive distribution requires the following steps. First, an independent sample from the BNN posterior is needed, which costs $O(T(NMD+BD))$, where $B$ is the number of minibatch points used in the likelihood (since we sample with SGLD). Then once an independent sample is identified, we evaluate it on the test points, which costs $O(PD)$. In total, generating an independent sample from the posterior predictive distribution costs $O(T(NMD+BD)+PD)$. If we were to use a naive GP, we need to generate a sample from the GP prior on the minibatch points in order to evaluate the likelihood, and then also sample the posterior predictive distribution. Recalling that sampling a GP scales with the cube (for the Cholesky decomposition required to identify the matrix inverse), this costs in total $O(T(1/3B^{3}+B)+1/3P^{3})$, where again, we assume the posterior is sampled with SGLD to exploit minibatching. For a more scalable GP method, we also consider KISS-GP, which scales linearly in the number of minibatch points and quadratically in the number of test points [61], leading to a cost of $O(2TB+B^{2})$. As there are many hyperparameters to tune, we consider a high-cost and low-cost scenario, with the values for each reported in Table 1.
In Fig. 1, we compare the computational cost of generating a single sample from the posterior predictive distribution for each method. We mark where each method crosses the threshold beyond the capabilities of a single NVIDIA H100 GPU, which has a reported computational capability of up to $989$ teraFLOPs. Even in the high-cost case, the BNN with the Mercer prior is able to handle posteriors beyond a billion test points. In the low-cost case, the Mercer prior outperforms the other methods even at the relatively low number of test points close to $100$.
@@ 4 Case study: drawing neural networks from Brownian motion
for $n\in\mathbb{N}$. For the field parameterization, we let $u_{\theta}(t)=tf(t;\theta)$, where $f$ is a single-layer neural network with the sigmoid activation function $\sigma(t)=(1+e^{-t})^{-1}$. We use this specific parameterization to automatically satisfy the condition that $u_{\theta}(0)=0$ in order to follow a Brownian motion. Throughout the experiments, we vary both the width of the network and the number of kernel terms $K$. In addition to this, we add a Fourier feature input layer to the networks [52]. This is in effort to combat the spectral bias of neural networks and allows the parameterization to more easily capture the characteristic high-frequency nature of Brownian motion. The inclusion of the Fourier feature layer helps SGLD identify the high-probability region faster. In principle, this could be omitted, although this would need significantly more samples of SGLD to truly characterize the posterior. We use the open source package Equinox [25], which is built on top of JAX [8], for implementation, and for sampling with SGLD, we use the samplers found in the Blackjax [9] library.
@@
In Fig. 3, we compare the true covariance function of Brownian motion $k(s,t)=\min(s,t)$ to the empirical covariance matrix of the BNN samples. When calculating the empirical covariance, we take $N_{F}=100,000$ samples, with each network being evaluated on $m=1,000$ points. The matrix of absolute errors for this case is shown in Fig. 4(a). We report a maximum error of $\approx 0.43$, which is located at the tip $s=t=1$. This represents a discrepancy of less than $5\%$. Recall that (centered) GPs are uniquely determined by their covariance functions. This level of agreement between the covariances of true Brownian motion and our BNNs therefore suggests that the BNNs are in fact approximately sampled from the Wiener measure.
@@
The main difference between the BNNs and true Brownian motion is that the BNNs still have smoothness to them. We can observe this in the covariance functions, as the BNN covariance is a bit more blurred. This fact should not be a surprise, as the finite-width neural network will be infinitely differentiable for any set of parameters $\theta$ since we have chosen the sigmoid activation function, while Brownian motion is famously nowhere differentiable. One may therefore be tempted to replace the activation function with one which produces non-differentiable neural networks, such as the rectified linear unit (ReLU). However, doing so will lead to an ill-defined prior, even for a finite-width neural network. Recall that before building an unbiased estimate, the BNN prior looks like $p(\theta)\propto\exp\left(-1/2\langle u_{\theta},S^{-1}u_{\theta}\rangle\right)$, where we then express $S^{-1}$ in terms of its Mercer representation. For Brownian motion, one can show that this prior also looks like

### v2

@@ 3 The Mercer prior
Proposition 3.1 provides an unbiased estimator of the Mercer prior as given by eq. (3.4). The estimate exploits the available spectral representation of the covariance operator $S$ for the measure we wish to draw the BNN from. In this way, we avoid ever needing to compute the integrals explicitly, identify the inverse covariance kernel analytically, or invert a large covariance matrix as is typical when sampling from a GP. At any given step, to compute the estimate, we simply need to minibatch points in the domain in order to approximate the integrals and select random eigenvalues/eigenfunctions associated with $S$. The properties of SGLD guarantee convergence to the correct probability distribution even for all batch sizes equal to one, showing the potential for massive parallelization. For convenience, we summarize a sampling scheme for the Mercer prior in Algorithm 1.
In proposition 3.1, we have chosen to minibatch points in $\Omega$ uniformly, which does not allow the domain to be infinite. This choice was made in order to simplify the method and it is not a strict requirement. If one wished to expand the domain to $\mathbb{R}^{d}$, importance sampling can be used to turn the integrals $\langle u_{\theta},\phi_{n}\rangle$ into expectations. Specifically, if we instead sample $t_{\beta}$ and $t_{\gamma}$ from distributions supported on $\mathbb{R}^{d}$, $q(t)$ and $w(t)$, respectively, we approximate each inner product using
$\langle u_{\theta},\phi_{n}\rangle\approx\sum_{\beta=1}^{M_{1}}\frac{u_{\theta}(t_{\beta})\phi_{n}(t_{\beta})}{q(t_{\beta})},\quad$
@@
1:initial parameters $\theta_{0}$, total iterations $T$, spectral indices NN, spatial minibatches $M_{1}$, $M_{2}$, initial learning rate $\varepsilon_{0}$
2:Parameter samples $\{\theta_{j}\}_{j=0}^{T}$
3:initialize parameters $\theta_{j}\leftarrow\theta_{0}$ and learning rate $\varepsilon_{j}\leftarrow\varepsilon_{0}$;
4:for $j=0$ to T1T-1 do
  sample minibatch $\mathcal{B}_{1}=\{t_{\beta}\}_{\beta=1}^{M_{1}}\sim U(\Omega)$;
  sample minibatch $\mathcal{B}_{2}=\{t_{\gamma}\}_{\gamma=1}^{M_{2}}\sim U(\Omega)$;
  sample spectral index minibatch $\mathcal{I}=\{n_{\alpha}\}_{\alpha=1}^{N}\sim p(n)$;
  sample SGLD noise $\eta_{j}\sim\mathcal{N}(0,\varepsilon_{j}I)$;
 compute inner product estimators:
$(v_{1})_{\alpha}\leftarrow\sum_{t\in\mathcal{B}_{1}}u_{\theta_{j}}(x)\,\phi_{n_{\alpha}}(t),\qquad n_{\alpha}\in\mathcal{I}$
$(v_{2})_{\alpha}\leftarrow\sum_{t\in\mathcal{B}_{2}}u_{\theta_{j}}(t)\,\phi_{n_{\alpha}}(t),\qquad n_{\alpha}\in\mathcal{I}\;$
 form unbiased Monte Carlo estimator:
$\mathcal{S}_{N,M_{1},M_{2}}(\theta_{j})\leftarrow-\frac{1}{2}\frac{|\Omega|^{2}}{NM_{1}M_{2}}\sum_{n_{\alpha}\in\mathcal{I}}\lambda_{n_{\alpha}}^{-1}p(n_{\alpha})^{-1}(v_{1})_{\alpha}(v_{2})_{\alpha}\;$
 SGLD update:
$\theta_{j+1}\leftarrow\theta_{j}+\frac{1}{2}\varepsilon_{j}\nabla_{\theta}\mathcal{S}_{N,M_{1},M_{2}}(\theta_{j})+\eta_{j}\;$
 update learning rate $\varepsilon_{j+1}$;
5:return parameter samples $\{\theta_{j}\}_{j=0}^{T}$
Algorithm 1 Mercer prior sampler
@@ 3.1 A cruder approximation
$\mathcal{S}_{N,M_{1},M_{2}}=-\frac{1}{2}\frac{K|\Omega|^{2}}{NM_{1}M_{2}}\sum_{\alpha=1}^{N}\frac{1}{\lambda_{n_{\alpha}}}\left\{\sum_{\beta=1}^{M_{1}}u_{\theta}(t_{\beta})\phi_{n_{\alpha}}(t_{\beta})\right\}\left\{\sum_{\gamma=1}^{M_{2}}u_{\theta}(t_{\gamma})\phi_{n_{\alpha}}(t_{\gamma})\right\},$ (3.7)
@@ 3.2 Specifying Gaussian process covariance functions from the ground up
Observe that the method builds sampling approximations of GPs using only the eigenvalues and eigenfunctions of the covariance kernel. This means that the covariance kernel does not need to be explicitly written. The benefit of this is that we may specify the desired GP completely through the eigenvalues and eigenfunctions, which allows us to construct BNNs for specialized tasks. We describe two possible approaches for constructing a desired kernel.
@@
$k(s,t)=\sum_{n=1}^{\infty}\lambda_{n}\phi_{n}(s)\phi_{n}(t)$ (3.8)
defines a valid covariance kernel for a GP. Note that this also holds if the series is truncated at some finite number of terms. In this way, we may start by choosing the eigenvalues and eigenfunctions so that the GP follows some desired behavior, e.g., smoothness, which is then transferred into the behavior of the BNN. The importance of this is that for GP kernels commonly used in machine learning tasks, the eigenvalues and eigenfunctions are currently not known or not easily computed. As an example, the eigenvalues and eigenfunctions of the squared exponential kernel $k(s,t)=\exp(-\|s-t\|^{2}/\ell^{2})$, which is perhaps the most commonly used kernel, are only available under the Gaussian measure $\mathcal{N}(0,\sigma^{2})$, for some variance $\sigma^{2}>0$ [68]. The eigenfunctions are then defined via transformations of the Hermite polynomials. In numerical experiments not covered in this work, we found that we could not include enough of these eigenfunctions in the expansion to build a good approximation of the GP. This was because after roughly $100$ terms, the computations became very unstable, mostly attributed to the fact that the normalization constants became exceptionally large.
@@
    For all fixed $t\in\Omega$, $k(\cdot,t)\in H_{k}$.
@@
    For all fixed $t\in\Omega$, $u(t)=\langle u,k(\cdot,t)\rangle_{H_{k}}$ for each $u\in H_{k}$.
@@
The second method involves damping an existing covariance with a Mercer representation. One specific target application we have in mind for the Mercer prior is in infinite-dimensional Bayesian inverse problems, where one would look to replace the usual prior Gaussian measure with the corresponding BNN for scalability. In such inverse problems, one often works with a ‘Laplace-like’ covariance, see [51]. This is first described by the baseline Gaussian measure $\mathcal{N}(0,\Delta^{-1})$, where $\Delta$ is the Laplacian with Dirichlet boundary conditions. This measure is famously known as the Brownian bridge measure. On $[0,1]$, the kernel of the inverse Laplacian operator $\Delta^{-1}$ is simply $k(s,t)=\min\{s,t\}-st$. On the $d$-dimensional unit cube, the kernel is best expressed in terms of its eigenvalues and eigenfunctions, given by
$\lambda_{n_{1},\dots,n_{d}}$ $=\frac{1}{(n_{1}\cdots n_{d})^{2}\pi^{2}}$
$\phi_{n_{1},\dots,n_{d}}(t)$ $=2^{d/2}\sin(n_{1}\pi t)\cdots\sin(n_{d}\pi t),$
respectively [42, Sec. 8.2.2-16]. Hence, by Mercer representation the kernel can be written explicitly as
$k(s,t)=2^{d}\sum_{n_{1}=1}^{\infty}\cdots\sum_{n_{d}=1}^{\infty}\frac{\sin(n_{1}\pi s)\cdots\sin(n_{d}\pi s)\sin(n_{1}\pi t)\cdots\sin(n_{d}\pi t)}{(n_{1}\cdots n_{d})^{2}\pi^{2}}.$
@@ 3.3 Computational cost analysis
$\mathcal{S}_{M_{1}}=\sum_{\beta=1}^{M_{1}}u_{\theta}(t_{\beta})\phi_{n_{\alpha}}(t_{\beta}),$
and similarly for $M_{2}$. For simplicity, we will assume that the eigenvalues and eigenfunctions are known analytically and do not need to be numerically computed, e.g., with a Nyström approximation [60]. This holds for the wide class of kernels based on the Laplacian operator commonly used in Bayesian inverse problems, or when engineering a new kernel, both discussed in Sec. 3.2.
Now, each term $u_{\theta}(t)$ requires a forward pass through a neural network, which costs $O(D)$. Evaluating an eigenfunction $\phi_{n_{\alpha}}(t)$ costs $O(1)$, unless the eigenfunction is a special function, e.g. a Bessel function, which we will ignore. So, this incurs a cost of $O((M_{1}+M_{2})D)=O(MD)$ per eigen-index if we take $M_{1}=M_{2}=M$. Multiplying by $N$ gives the total leading cost $O(NMD)$. To be completely fair, we acknowledge the fact that SGLD tends to mix very slowly, so the unbiased estimate needs to be computed many times in order to generate one independent posterior draw. If we let $T$ denote the number of SGLD steps needed for an independent sample, the total cost is $O(TNMD)$. The mixing rate can be improved by implementing a variant of SGLD, such as preconditioning [58] or cyclical SGLD [67]. Alternatively, the prior can be characterized by other methods which are known to work with unbiased estimates, such as Hamiltonian Monte Carlo with energy conserving subsampling [12], or with variational inference [6].
@@
If we replace the GP with a BNN equipped with the Mercer prior, generating a sample from the posterior predictive distribution requires the following steps. First, an independent sample from the BNN posterior is needed, which costs $O(T(NMD+BD))$, where $B$ is the number of minibatch points used in the likelihood (since we sample with SGLD). Then once an independent sample is identified, we evaluate it on the test points, which costs $O(PD)$. In total, generating an independent sample from the posterior predictive distribution costs $O(T(NMD+BD)+PD)$. If we were to use a naive GP, we need to generate a sample from the GP prior on the minibatch points in order to evaluate the likelihood, and then also sample the posterior predictive distribution. Sampling a GP scales with the cube (for the Cholesky decomposition required to identify the matrix inverse), so this costs in total $O(T(1/3B^{3}+B)+1/3P^{3})$, where again, we assume the posterior is sampled with SGLD to exploit minibatching. For a more scalable GP method, we also consider KISS-GP, which scales linearly in the number of minibatch points and quadratically in the number of test points [64], leading to a cost of $O(2TB+B^{2})$. As there are many hyperparameters to tune, we consider a high cost and low cost scenario, with the values for each reported in Table 1.
In Fig. 1, we compare the computational cost of generating a single sample from the posterior predictive distribution for each method. We mark where each method crosses the threshold beyond the capabilities of a single NVIDIA H100 GPU, which has a reported computational capability of up to $989$ teraFLOPs. Even in the high cost case, the BNN with the Mercer prior is able to handle posteriors beyond a billion test points. In the low cost case, the Mercer prior outperforms the other methods even at the relatively low number of test points close to $100$.
@@
3.4 Proposed directions for hyperparameter selection
which defines the properly normalized prior $p(\theta|\lambda)=Z(\lambda)^{-1}\exp(-E(\theta;\lambda))$. To connect the technique back to its inspiration in field theories, $E$ can be viewed as the energy of the system, and $Z(\lambda)$ is the corresponding partition function. Under this view, the prior assigns higher probability density to neural networks which are closer to the minimum energy state.
A fully Bayesian treatment over both $\theta$ and $\lambda$ requires accounting for the variation of $Z(\lambda)$, as the joint posterior would be
once one picks a hyperprior $p(\lambda)$. Of course, this is unlikely to be tractable due to the complicated map defined by the neural network architecture within $Z(\lambda)$. Ignoring this subtlety and dropping the partition function introduces bias into the inference whenever $\lambda$ is learned jointly with $\theta$. In our examples in this work, we will keep $\lambda$ fixed so that $Z(\lambda)$ is simply a constant and can be dropped when sampling with SGLD. Our primary concern is how the Mercer prior performs as an approximation to GPs in applications, not necessarily on how any hyperparameters are learned.
If desired, one approach to train any hyperparameters would be to work with the marginal $p(\lambda|y)=\int_{\Theta}p(\theta,\lambda|y)d\theta$, which can be characterized through the following identity
$\nabla_{\lambda}\log p(\lambda|y)=\mathbb{E}_{p(\theta|y,\lambda)}[-\nabla_{\lambda}E(\theta;\lambda)]-\mathbb{E}_{p(\theta|\lambda)}[-\nabla_{\lambda}E(\theta;\lambda)]-\nabla_{\lambda}\log p(\lambda),$ (3.9)
which is derived in [3]. Let us consider the simplified case where $p(\lambda)$ is flat, so that only the expectations appear. Then, the practice reduces to model selection by marginal likelihood [59]. This expression states that the gradient of the $\log$-evidence with respect to the hyperparameters is the difference between the expected energy gradients under the posterior and prior, respectively. Intuitively, the first term dictates how the energy landscape shifts under the posterior, i.e., it points towards the direction of $\lambda$ which reduces the energy. The second term is a correction coming from the change in normalization constant. It is a baseline shift in energy that happens even without data. Again to connect back to physics, the derivative of the energy $\nabla_{\lambda}E(\theta;\lambda)$ can be viewed as a force, so eq. (3.9) describes the expected excess force that $\lambda$ is subjected to induced by the data. Probabilistically, this is exactly the score function gradient estimator for the likelihood $\nabla_{\lambda}\mathbb{E}_{p(\theta|\lambda)}[p(y|\theta)]=\mathbb{E}_{p(\theta|\lambda)}[p(y|\theta)\nabla_{\lambda}\log p(\theta|\lambda)]$ [62]. This can be seen by first observing that
by using the property of the partition function that $\nabla_{\lambda}\log Z(\lambda)=-\mathbb{E}_{p(\theta|\lambda)}[\nabla_{\lambda}E(\theta;\lambda)]$. Taking the posterior expectation reveals exactly the form of eq. (3.9).
In practice, the identity eq. (3.9) enables empirical Bayes optimization of $\lambda$ using Monte Carlo estimates of the two expectations, which can be obtained via SGLD samples from the posterior and prior with fixed hyperparameters. A fully Bayesian approach would be possible with nested sampling as found in [3]. This approach uses alternating SGLD updates for $\theta$ to approximate the expectations, which then inform an SGLD step for $\lambda$ in order to sample the joint posterior without explicitly computing $Z(\lambda)$. While exact, methods that require evaluating or estimating $Z(\lambda)$, such as pseudo-marginal Metropolis-Hastings, are typically computationally demanding. One could also optimize $\lambda$ using the score function gradient estimator, e.g., with Adam [26].
@@ 4 Case study: drawing neural networks from Brownian motion
for $n\in\mathbb{N}$. For the field parameterization, we let $u_{\theta}(t)=tf(t;\theta)$, where $f$ is a single-layer neural network with the sigmoid activation function $\sigma(t)=(1+e^{-t})^{-1}$. We use this specific parameterization to automatically satisfy the condition that $u_{\theta}(0)=0$ in order to follow a Brownian motion. This ensures that each BNN sample belongs to the space $C_{0}([0,1])$. Throughout the experiments, we vary both the width of the network and the number of kernel terms $K$. In addition to this, we add a Fourier feature input layer to the networks [53]. This is in effort to combat the spectral bias of neural networks and allows the parameterization to more easily capture the characteristic high-frequency nature of Brownian motion. The inclusion of the Fourier feature layer helps SGLD identify the high-probability region faster. In principle, this could be omitted, although this would need significantly more samples of SGLD to truly characterize the posterior. We use the open source package Equinox [25], which is built on top of JAX [8], for implementation, and for sampling with SGLD, we use the samplers found in the Blackjax [9] library.
@@
In Fig. 3, we compare the true covariance function of Brownian motion $k(s,t)=\min(s,t)$ to the empirical covariance matrix of the BNN samples. When calculating the empirical covariance, we take $N_{F}=100,000$ samples, with each network being evaluated on $m=1,000$ points. We then perform two statistical tests to qualitatively assess the performance of the method. First, we directly compare the true covariance of Brownian motion to the empirical covariance obtained by the Mercer prior. The matrix of absolute errors for this case is shown in Fig. 4(a). We report a maximum error of $\approx 0.43$, which is located at the tip $s=t=1$. This represents a discrepancy of less than $5\%$. Recall that (centered) GPs are uniquely determined by their covariance functions. This level of agreement between the covariances of true Brownian motion and our BNNs therefore suggests that the BNNs are in fact approximately sampled from the Wiener measure.
@@
The main difference between the BNNs and true Brownian motion is that the BNNs still have smoothness to them. We can observe this in the covariance functions, as the BNN covariance is a bit more blurred. This can also be seen by the fact that most of the error in the covariance concentrates along the center line $s=t$, which suggests the BNNs have difficultly in capturing the extreme frequencies. This fact should not be a surprise, as the finite-width neural network will be infinitely differentiable for any set of parameters $\theta$ since we have chosen the sigmoid activation function, while Brownian motion is famously nowhere differentiable. One may therefore be tempted to replace the architecture with one which produces non-differentiable neural networks. However, doing so will lead to an ill-defined prior, even for a finite-width neural network. Recall that before building an unbiased estimate, the BNN prior looks like $p(\theta)\propto\exp\left(-1/2\langle u_{\theta},S^{-1}u_{\theta}\rangle\right)$, where we then express $S^{-1}$ in terms of its Mercer representation. For Brownian motion, one can show that this prior also looks like